TIP: Click on subject to list as thread! ANSI
echo: sb-nasa_news
to: All
from: Hugh S. Gregory
date: 2003-02-16 23:12:00
subject: 1\30 Pt-1 ISS On Orbit Status 30-01-2003

This Echo is READ ONLY !   NO Un-Authorized Messages Please!
 ~~~~~~~~~~~~~~~~~~~~~~~~   ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
1\30 ISS On-Orbit Status 30-01-2003
Part 1 of 2

ISS On-Orbit Status 1/30/03

All ISS systems continue to function nominally, except as noted 
previously or below.

Wake-up was at the regular 1:00am EST, followed by post-sleep 
activities and breakfast (begin of sleep time: 4:30pm).

FE-1 Nikolai Budarin and CDR Kenneth Bowersox worked on the Progress 
9P spacecraft, installing the docking mechanism (StM) of the cargo 
vehicle's SSVP docking and internal transfer system. [The SSVP on 
Progress (and Soyuz) is of the "probe and cone" (
"classic") type, to 
distinguish it from the structurally more robust hybrid ASP type used 
on ISS for connecting Russian modules together. Besides acting as 
structural and mechanical connection allowing passage of crew and 
cargo, it also has connections for power, data (command and control), 
gases and fluids (propellants).]

Afterwards, Budarin deinstalled and removed no-longer-needed 
electronic equipment from the Progress, viz., a local signal switch 
device (LKT/TA251) of the BITS onboard telemetry system, with its ROM 
memory (PZU), for return to Earth and recycling on a future Progress.

The crew continued stowing disposable cargo on Progress, while 
video-taping the activities with the U.S. Sony PD-100 DVCAM. The 
recording was then downlinked via Ku- and S-band at 7:10am.

Budarin conducted his regular daily maintenance checkup of the Russian 
BIO-5 Rasteniya-2/Lada-2 ("Plants-2") plant growth experiment. Later, 
he copied its accumulated photo/data files to a floppy disk for 
transfer to the Russian Laptop 3 and subsequent downlink via 
Regul-Packet.

FE-2/SO Don Pettit terminated the regeneration cycle on the first EMU 
Metox (metal oxide) CO2 filter in the Airlock's bake-out oven, then 
started it on the second canister.

Early in the morning, Pettit took carbon dioxide partial pressure 
(ppCO2) readings of the cabin air in Lab and SM with the portable ACS 
CDMK (atmosphere control system/CO2 monitoring kit). A second data 
take is scheduled for shortly before sleep time. [ppCO2 levels have 
shown a gradual increase, going above the lower limit of 5.3 mmHg over 
the last weekend. The crew has been instructed to resume daily CDMK 
readings (morning and evening) for the near-term, while specialists 
investigate the cause. The Service Module (SM) CO2 sensor is reading 
ppCO2 levels well below the Lab MCA (major constituent analyzer) and 
CDMK measurements (see Systems Status, below).]

Air samples were collected by Pettit in the Lab, SM, and Node with the 
MAS (microbial air sampler) kit for bacterial and fungal analysis, to 
be performed on Days 2 and 5 of incubation in Petri dishes. The 
collected samples were photographed by Bowersox for the record.

Sox then conducted surface sample swabs in the Lab, SM and Node with 
the SSK (surface sample kit), for cultivation. The colony growth on 
the sampling slides will be analyzed after two and five days of 
incubation. The activities and SSK samples were photographed by Don 
for the record

Troubleshooting continues on the ER3 ARIS (Express rack 3/active rack
isolation system), after the crew was unable to perform the pushrod
calibration on 1/28. [Under investigation is a possible issue with the 
new software load (file transfer from the ER3 memory unit {EMU} to the 
ARIS controller). To protect the ARIS hardware, the ARIS alignment 
guides need to remain installed until the problem is fixed. Some crew 
support may be requested for tomorrow (Friday).]

Early in the morning, Don Pettit powered up the HRF (human research
facility) for checking out the FOOT (foot/ground reaction forces 
during space flight) experiment's montage file, which he then copied 
for downlink.  The HRF computer's data files were subsequently 
downlinked on ground command, and the HRF was deactivated again.

Pettit also worked on the EarthKAM (EK), changing its optics to the 
180-mm lens. Unattended (ground-commanded) uplink and downlink of EK 
payload control and data files through OCA took place throughout the 
day. [The EK team, on behalf of the 29 participating middle schools 
across the USA, Mexico and Japan, sent up sincere thanks for the 
crew's support in setting up EK on the OpsLAN (operations/local area 
network). So far, the ground has commanded and received over 280 
spectacular images of the Earth, and two favorite ones were uplinked 
for the crew's perusal. EarthKAM images can be found at
http://datasystem.earthkam..ucsd.edu/cgi-bin/datasys/ek_images_station 
].

Budarin completed the regular daily task of SOSh life support systems
maintenance, and performed the periodic routine inspection of the BRPK
air/condensate water separator in the SM, while Pettit took care of 
the daily IMS database update file preparation.

A Lab RPCM (remote power controller module), which powers the VTR1 
(video tape recorder #1), has tripped open. VTR1 was declared "off 
limits" until the problem is resolved. [Telemetry revealed no evidence 
ofa VTRovercurrent that would have tripped the fuse. Although this 
could simply be a nuisance trip, the signature is similar to a problem 
seen in the past, which eventually led to a POR (power-on reset). This 
would require the ground to run a specific nominal procedure, with the 
VTR off.]

Early in the morning, MCC-H and its support group in Moscow (HSG) 
performed a BCC (backup control center) dry-run in test mode, with no 
involvement of ISS crew or vehicle. [In test mode, HSG (as BCC) was to 
send a single not-for-uplink PPCP (preplanned command packet) to 
MCC-M/TsUP, then to cancel it after successful transmission. MCC-H was 
to play back the ACT (American contingency telemetry) recorded during 
the Hurricane Lili contingency operations last year (when MCC-H was 
shut down at 4:00am EDT on 10/2), with HSG confirming successful 
receipt. TsUP commanding to the Russian segment (RS) via U.S. S-band, 
as well as MCC-H-to-TsUP flow of status telemetry were also not 
affected.]

Another water venting from the Lab condensate collection tank is 
tentatively scheduled for Friday next week (2/7), when tank contents 
are predicted to reach the upper limit of 100 lbs. During the dump, 
the Lab window shutter will be closed by the crew to protect it 
against contamination. [According to applicable Flight Rule, the 
condensate tank should preferably be off-loaded into a CWC 
(contingency water container) for processing/recycling in theRS. 
Currently, however, US condensate can not be processed until about 
April this year due to previous failures of the Russian SRV-K water 
processor system. The excess water must therefore be vented overboard 
(dual-stream, to cancel out propulsive reaction on the vehicle), in 
LVLH (+XVV) attitude.]

(continued)

---
* Origin: SpaceBase[tm] Vancouver Canada [3 Lines] 604-473-9357 (1:153/719)
SEEN-BY: 633/267 270
@PATH: 153/719 715 7715 140/1 106/2000 633/267

SOURCE: echomail via fidonet.ozzmosis.com

Email questions or comments to sysop@ipingthereforeiam.com
All parts of this website painstakingly hand-crafted in the U.S.A.!
IPTIA BBS/MUD/Terminal/Game Server List, © 2025 IPTIA Consulting™.